OpenStudy (anonymous):

which is a power of 10? 10-3 1^100 5,000 100.01

OpenStudy (anonymous):

100.01 is not a power of 10 because it is something like 10001 with some power 10^1=10, 10^2=100 5 is not the power of a whole number, either

OpenStudy (anonymous):

10-3 = 7 ?? 7 is not a power of 10

OpenStudy (anonymous):

the answer 10-3

OpenStudy (anonymous):

right ?

OpenStudy (anonymous):

10-3 = 7 7 is not a power of 10

OpenStudy (anonymous):

10^-3 : power of 10 10-3 : not a power of 10!

OpenStudy (anonymous):

yes 10^-3

OpenStudy (anonymous):

oh, okay! yes that's a power of 10

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

1^100 = 1 10^0 = 1 maybe they also want that one, it's not written as a power of 10 but it does have the same value.....

OpenStudy (anonymous):

so what is the answer 1^100 or 10^-3??

OpenStudy (anonymous):

I'd say 10^-3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Np